# Batalin-Fradkin-Vilkovisky Quantization of FLPR model

### v1

@@ 3 FLPR model in Cartesian coordinates: Constraint structure
where $\dot{x}$, $\dot{y}$ and $\dot{z}$ represent the generalized velocities, $\zeta$ denotes the gauge variable and $g$ is a coupling constant. The canonical momenta corresponding to the coordinates $x$, $y$, $z$ and $\zeta$ are obtained as follows:
$p_{x}=\dot{x}+gy\zeta,\qquad p_{y}=\dot{y}-gx\zeta,\qquad p_{z}=\dot{z}-\zeta,\qquad p_{\zeta}=0.$ (20)
@@
There are only two constraints $(\tilde{\Omega}_{1}$ and $\tilde{\Omega}_{2})$ in the theory and these are first-class in nature [1, 17].
@@ 3.1 BFV quantization
To employ BFV formalism, we introduce a pair of canonically conjugate ghost and anti-ghost variables along with their auxiliary variables as [7, 10]
$\big{(}{\cal{C}}^{i},{\cal\bar{P}}_{i}\big{)},\qquad\big{(}{\cal{P}}^{i},{\cal\bar{C}}_{i}\big{)},\qquad\big{(}{N}^{i},{B}_{i}\big{)},$ (23)
where $i=1,2.$ These variables satisfy the super-Poisson algebra given in Eq. (5). As per the BFV formulation, the nilpotent BRST charge ($\tilde{Q}$) and the fermionic gauge-fixing function ($\tilde{\Phi}$) are defined using constraints and gauge-fixing conditions, respectively, as follows:
@@
$\tilde{H}_{m}$ $=$ $\tilde{H}_{c}+\bar{{\cal{P}}_{a}}V^{a}_{b}{\cal{C}}^{b},$ (26)
$=$ $\frac{1}{2}\Big{(}p_{x}^{2}+p_{y}^{2}+p_{z}^{2}\Big{)}+\zeta\Big{(}g(xp_{y}-yp_{x})+p_{z}\Big{)}+U(x^{2}+y^{2})+{\cal\bar{P}}_{2}\,{\cal{C}}^{1},$
where we assume that the constraints fulfill the constraint algebra $\{H_{c},\Omega_{b}\}=\Omega_{a}V^{a}_{b}$. Moreover, $\tilde{Q}$, $\tilde{\Phi}$ and $\tilde{H}_{m}$ satisfy the algebra as listed in Eqs. (10) and (11).
Hence, the effective action in the extended phase space can be given as
$\tilde{S}_{eff}=\int\text{d}t\Big{[}p_{x}\dot{x}+p_{y}\dot{y}+p_{z}\dot{z}+p_{\zeta}\dot{\zeta}+B_{2}\dot{N^{2}}+{\cal\bar{P}}_{1}{\cal{\dot{C}}}^{1}+{\cal\bar{P}}_{2}{\cal{\dot{C}}}^{2}+{\cal\bar{C}}_{2}{\cal{\dot{P}}}^{2}\Big{]}-\tilde{H}_{total},$ (27)
where $\tilde{H}_{total}=\tilde{H}_{m}+\{\tilde{Q},\tilde{\Phi}\}$. In the above, the terms $\int dt\,(B_{1}\,\dot{N^{1}}+{\cal\bar{C}}_{1}\,{\cal{\dot{P}}}^{1}$ ) =  $\{\tilde{Q},\int dt\,\;{\cal\bar{C}}_{1}\,\dot{N}^{1}\}$ in Legendre transformations are suppressed by replacing $\tilde{\chi}^{1}$ with $\tilde{\chi}^{1}+\dot{N}^{1}$. Now, in order to obtain the effective action, first we perform the path integration over the variables $B_{1}$, $N^{1}$, ${\cal\bar{C}}_{1}$, ${\cal{P}}^{1}$, ${\cal\bar{P}}_{1}$, ${\cal{C}}^{1}$, $\zeta$ and $p_{\zeta}$ using Gaussian integration. Thus, we obtain
$\tilde{S}_{eff}$ $=$ $\int\text{d}t\Big{[}p_{x}\dot{x}+p_{y}\dot{y}+p_{z}\dot{z}+B\dot{N}+{\cal\bar{P}}{\cal{\dot{C}}}+{\cal\bar{C}}{\cal{\dot{P}}}-\frac{1}{2}p_{x}^{2}-\frac{1}{2}p_{y}^{2}-\frac{1}{2}p_{z}^{2}-U(x^{2}+y^{2})$ (28)
@@
here we have redefined $N^{2}\equiv N$, $B_{2}\equiv B$, ${\cal\bar{C}}_{2}\equiv{\cal\bar{C}}$, ${\cal{C}}^{2}\equiv{\cal{C}}$, ${\cal\bar{P}}_{2}\equiv{\cal\bar{P}}$ and ${\cal{P}}^{2}\equiv{\cal{P}}$. Second, taking the variations over $p_{x}$, $p_{y}$, $p_{z}$, $\cal{P}$} and $\cal\bar{P}$} yields the following relations
@@
$\tilde{S}_{eff}$ $=$ $\int\text{d}t\,\Big{[}\frac{1}{2}\big{(}(\dot{x}+gy\zeta)^{2}+(\dot{y}-gx\zeta)^{2}+(\dot{z}-\zeta)^{2}\big{)}+U(x^{2}+y^{2})$ (30)
@@
$\begin{split}&\tilde{\delta}_{b}x=-gy{\cal{C}},\qquad\tilde{\delta}_{b}y=gx{\cal{C}},\qquad\tilde{\delta}_{b}z={\cal{C}},\qquad\tilde{\delta}_{b}\zeta=\dot{\cal{C}},\\ &\tilde{\delta}_{b}{\cal{C}}=0,\qquad\tilde{\delta}_{b}\bar{\cal{C}}=-B,\qquad\tilde{\delta}_{b}B=0,\end{split}$ (31)
@@
$\tilde{S}_{eff}$ $=$ $\int\text{d}t\,\Big{[}\frac{1}{2}\big{(}(\dot{x}+gy\zeta)^{2}+(\dot{y}-gx\zeta)^{2}+(\dot{z}-\zeta)^{2}\big{)}+U(x^{2}+y^{2})$ (32)
@@ 4 Physicality criteria: BRST charge
are the generators of BRST symmetry transformations (cf. (17) and (31)) under which the effective actions $(S_{eff}$ and $\tilde{S}_{eff})$ remain invariant, respectively. In the above equation, we have used the Euler-Lagrange equations of motion ${\cal{P}}=\dot{\cal{C}}$ and the expression for momenta $p_{\zeta}=B$. These charges are also nilpotent of order two (cf. (10)).
@@
It is important to note that $(gp_{\theta}+p_{z})\approx 0$ and $p_{\zeta}\approx 0$ are first-class constraints of the system in the polar coordinates whereas $\big{(}g\,(xp_{y}-yp_{x})+p_{z}\big{)}\,\approx 0$ and $p_{\zeta}\,\approx 0$ are first-class constraints in Cartesian coordinates. Therefore, physical states are annihilated by the first-class constraints, which is consistent with the results in [17] and Dirac formalism.

### v2

@@ 4 FLPR model in Cartesian coordinates
where $\dot{x}$, $\dot{y}$ and $\dot{z}$ represent the generalized velocities, $\zeta$ denotes the gauge variable and $g$ is a coupling constant.
4.1 Constraint structure
The canonical momenta corresponding to the coordinates $x$, $y$, $z$ and $\zeta$ are obtained as follows:
$p_{x}=\dot{x}+gy\zeta,\quad p_{y}=\dot{y}-gx\zeta,\quad p_{z}=\dot{z}-\zeta,\quad p_{\zeta}=0.$ (30)
@@ 4.1 Constraint structure
There are only two constraints $(\tilde{\Omega}_{1}$ and $\tilde{\Omega}_{2})$ in the theory and these are first-class in nature (cf. [33] for details).
@@ 4.2 BFV quantization
As per the BFV formulation, the nilpotent BRST charge ($\tilde{Q}$) and the fermionic gauge-fixing function ($\tilde{\Phi}$) are defined using constraints and gauge-fixing conditions, respectively, as follows:
@@
$\tilde{H}_{m}=\frac{1}{2}\Big(p_{x}^{2}+p_{y}^{2}+p_{z}^{2}\Big)+\zeta\Big(g(xp_{y}-yp_{x})+p_{z}\Big)+U(x^{2}+y^{2})+{\cal\bar{P}}_{2}\,{\cal{C}}^{1}.$ (35)
Moreover, $\tilde{Q}$, $\tilde{\Phi}$ and $\tilde{H}_{m}$ satisfy the algebra as listed in Eqs. (11) and (12).
Hence, the generating functional in the extended phase space can be given as
$\tilde{\cal{Z}}=\int{\cal{D}}\tilde{\phi}\,\exp\big(i\,\tilde{S}_{eff}\big),$ (36)
where
${\cal{D}}\tilde{\phi}={\cal{D}}\big(x,p_{x},y,p_{y},z,p_{z},\zeta,p_{\zeta},N_{a},B_{a},{\cal{C}}^{a},\bar{{\cal{P}}}_{a},{\cal{P}}^{a},\bar{{\cal{C}}}_{a}\big)$
and
$\tilde{S}_{eff}=\int dt\Big[p_{x}\dot{x}+p_{y}\dot{y}+p_{z}\dot{z}+p_{\zeta}\dot{\zeta}+B_{a}\dot{N^{a}}+{\cal\bar{P}}_{a}{\cal{\dot{C}}}^{a}+{\cal\bar{C}}_{a}{\cal{\dot{P}}}^{a}-\tilde{H}_{total}\Big].$ (37)
Now, we suppress the terms $\int dt\,(B_{1}\,\dot{N^{1}}$+ ${\cal\bar{C}}_{1}\,{\cal{\dot{P}}}^{1}\,)\,$$\{\,\tilde{Q},\int dt\,\;{\cal\bar{C}}_{1}\,\dot{N}^{1}\}$ in the Legendre transformations by replacing $\tilde{\chi}^{1}$ with $\tilde{\chi}^{1}+\dot{N}^{1}$. Later, in order to obtain the effective action, first we perform the path integration over the variables $B_{1}$, $N^{1}$, ${\cal\bar{C}}_{1}$, ${\cal{P}}^{1}$, ${\cal\bar{P}}_{1}$, ${\cal{C}}^{1}$, $\zeta$ and $p_{\zeta}$ using Gaussian integration. Thus, we obtain
$\tilde{S}_{eff}$ $=$ $\int dt\Big[p_{x}\dot{x}+p_{y}\dot{y}+p_{z}\dot{z}+B\dot{N}+{\cal\bar{P}}{\cal{\dot{C}}}+{\cal\bar{C}}{\cal{\dot{P}}}-\frac{p_{x}^{2}}{2}-\frac{p_{y}^{2}}{2}-\frac{p_{z}^{2}}{2}-U(x^{2}+y^{2})$ (38)
@@
here, we have redefined $N^{2}\equiv N$, $B_{2}\equiv B$, ${\cal\bar{C}}_{2}\equiv{\cal\bar{C}}$, ${\cal{C}}^{2}\equiv{\cal{C}}$, ${\cal\bar{P}}_{2}\equiv{\cal\bar{P}}$ and ${\cal{P}}^{2}\equiv{\cal{P}}$. Second, taking the variations over $p_{x}$, $p_{y}$, $p_{z}$, $\cal{P}$} and $\cal\bar{P}$} yields the following relations
@@
$\tilde{S}_{eff}$ $=$ $\int dt\,\Big[\frac{1}{2}\,\Big((\dot{x}+gy\zeta)^{2}+(\dot{y}-gx\zeta)^{2}+(\dot{z}-\zeta)^{2}\Big)+U(x^{2}+y^{2})$ (40)
@@
$\tilde{s}_{b}x=-gy{\cal{C}},\quad\tilde{s}_{b}y=gx{\cal{C}},\quad\tilde{s}_{b}z={\cal{C}},\quad\tilde{s}_{b}\zeta=\dot{\cal{C}},\quad\tilde{s}_{b}{\cal{C}}=0,\quad\tilde{s}_{b}\bar{\cal{C}}=-B,\quad\tilde{s}_{b}B=0,$ (41)
@@
$\tilde{S}^{\prime}_{eff}$ $=$ $\int dt\,\Big[\,\frac{1}{2}\,\Big((\dot{x}+gy\zeta)^{2}+(\dot{y}-gx\zeta)^{2}+(\dot{z}-\zeta)^{2}\Big)+U(x^{2}+y^{2})$ (42)
@@ 5 Physicality criteria: BRST charge
are the generators of BRST symmetry transformations (cf. (27) and (41)) under which the effective actions $(S_{eff}$ and $\tilde{S}_{eff})$ remain quasi-invariant as
$s_{b}S_{eff}=\frac{d}{dt}(B{\dot{\cal{C}}}),\qquad\tilde{s}_{b}\tilde{S}_{eff}=\frac{d}{dt}(B{\dot{\cal{C}}}),$ (44)
respectively. It is worth mentioning that, in Eq. (43), we have used the Euler-Lagrange equations of motion ${\cal{P}}=\dot{\cal{C}}$ and the expression for momenta $p_{\zeta}=B$. Now, it is straightforward to prove the conservation of BRST charge $Q$ and $\tilde{Q}$ with the help of Euler-Lagrange equations of motion. These charges are also nilpotent of order two (cf. Eq. (11)) which signifies its fermionic character, since only operators with odd Grassmann parity can anti-commute with themselves to zero.
@@
It is important to note that $(gp_{\theta}+p_{z})\approx 0$ and $p_{\zeta}\approx 0$ are first-class constraints of the system in the polar coordinates, whereas $\big(g\,(xp_{y}-yp_{x})+p_{z}\big)\,\approx 0$ and $p_{\zeta}\,\approx 0$ are first-class constraints in Cartesian coordinates. Therefore, physical states are annihilated by first-class constraints, which is consistent with the results in [33] and the Dirac formalism.
6 Generalized BRST symmetry of FLPR model
We start with the partition function for the BFV-BRST invariant effective action of the FLPR model as
${\cal{Z}}=\int{\cal{D}}r\,{\cal{D}}\theta\,{\cal{D}}z\,{\cal{D}}\zeta\,{\cal{D}}B\,{\cal{D}}\bar{\cal{C}}\,{\cal{D}}{\cal{C}}\,e^{i\,S_{eff}},$ (48)
where the effective action $(S_{eff})$ is given by
$S_{eff}$ $=$ $\int dt\,\Big[\frac{1}{2}\,\dot{r}^{2}+\frac{1}{2}\,r^{2}\big(\dot{\theta}-g\zeta\big)^{2}+\frac{1}{2}\big(\dot{z}-\zeta\big)^{2}-V(r)$ (49)
$+$ $B\Big(\dot{\zeta}-z+\frac{1}{2}\,B\Big)-{\dot{\bar{\cal{C}}}}\,{\dot{\cal{C}}}-{\cal{\bar{C}}}\,{\cal{C}}\Big].$
This partition function and the effective action are invariant under the following off-shell nilpotent infinitesimal BRST transformations $(\delta_{b})$
${\delta}_{b}r=0,\quad{\delta}_{b}\theta=g{\cal{C}}\Lambda,\quad{\delta}_{b}z={\cal{C}}\Lambda,\quad{\delta}_{b}\zeta=\dot{\cal{C}}\Lambda,\quad{\delta}_{b}{\cal{C}}=0,\quad{\delta}_{b}{\cal\bar{C}}=-B\Lambda,\quad{\delta}_{b}B=0,$ (50)
where $\Lambda$ is an infinitesimal global parameter. As long as this parameter $\Lambda$ is anti-commuting and space-time independent, most of the characteristics of BRST transformations are unaffected by the field dependency or order of magnitude of the parameter $\Lambda$. However, we can generalize the BRST transformations to FFBRST transformations by making the anti-commuting infinitesimal parameter finite and field dependent. First, to make the parameter $\Lambda$ field dependent, we introduce a continuous arbitrary parameter $\kappa\,(0\leq\kappa\leq 1)$ such that
$\phi(t,\kappa=0)=\phi(t),\qquad\phi(t,\kappa=1)=\phi^{\prime}(t),$ (51)
where $\phi(t)$ denotes the generic variables of the theory. Now, the field-dependent infinitesimal BRST transformations can be written as
$\frac{d\theta(t,\kappa)}{d\kappa}=g{\cal{C}}\Theta^{\prime}[\phi(t,\kappa)],\quad\frac{dz(t,\kappa)}{d\kappa}={\cal{C}}\Theta^{\prime}[\phi(t,\kappa)],\quad\frac{d\zeta(t,\kappa)}{d\kappa}=\dot{\cal{C}}\Theta^{\prime}[\phi(t,\kappa)],$
$\frac{d\bar{\cal{C}}(t,\kappa)}{d\kappa}=-B\Theta^{\prime}[\phi(t,\kappa)],\quad\frac{dr(t,\kappa)}{d\kappa}=0,\quad\frac{d{\cal{C}}(t,\kappa)}{d\kappa}=0,\quad\frac{dB(t,\kappa)}{d\kappa}=0,$ (52)
where $\Theta^{\prime}[\phi(t,\kappa)]$ is a field dependent infinitesimal parameter. Then, we construct the FFBRST transformations $(\delta_{f})$ by integrating infinitesimal transformations in Eq. (6) from $\kappa=0$ to $\kappa=1$ as follows:
$\delta_{f}\,\theta=g\,{\cal{C}}\,\Theta[\phi(t)],\qquad\delta_{f}\,z={\cal C}\,\Theta[\phi(t)],\qquad\delta_{f}\zeta=\dot{\cal{C}}\,\Theta[\phi(t)],$
$\delta_{f}\bar{\cal{C}}=-B\,\Theta[\phi(t)],\qquad\delta_{f}\,r=0,\qquad\delta_{f}B=0,\qquad\delta_{f}{\cal{C}}=0,$ (53)
where the finite field-dependent parameter
$\Theta[\phi(t)]=\int^{1}_{0}d\kappa^{\prime}\Theta^{\prime}[\phi(t,\kappa^{\prime})].$ (54)
The effective action $(S_{eff})$ remains invariant under the FFBRST transformations (i.e., $\delta_{f}\,S_{eff}=0$). However, since the Grassmannian parameter is field-dependent, the path integral measure ${\prod_{\phi}}{\cal D}\phi(t)$ is not invariant under generalized BRST transformations. The Jacobian $J(\kappa)$ varies non-trivially and can be replaced by
$J(\kappa)\mapsto e^{-iS_{1}[\phi(t,\kappa),\kappa]},$ (55)
if and only if the condition
$\int\prod_{\phi}{\cal D}\phi(t,\kappa)\bigg[\dfrac{1}{J(\kappa)}\dfrac{dJ(\kappa)}{d\kappa}-i\,\dfrac{d}{d\kappa}S_{1}[\phi(t,\kappa),\kappa]\bigg]\,e^{i(S_{eff}+S_{1})}=0,$ (56)
is satisfied [16]. Here, $S_{1}[\phi(t,\kappa),\kappa]$ is some local functional of fields with boundary condition $S_{1}[\phi(t,\kappa),\kappa]_{\kappa=0}=0$ and
$\dfrac{1}{J(\kappa)}\dfrac{dJ(\kappa)}{d\kappa}=-\int dt\,\sum_{\phi}\bigg[\pm\,s_{b}\,\phi(t,\kappa)\dfrac{\Theta^{\prime}[\phi(t,\kappa)]}{\delta\phi(t,\kappa)}\bigg],$ (57)
where $\pm$ is the sign for bosonic and fermionic fields, respectively. Now, we construct the field-dependent BRST parameter as
$\Theta^{\prime}[\phi(t,\kappa)]=i\int dt\,\big(\dot{\zeta}(t,\kappa)-z(t,\kappa)\big)\,\bar{\cal C}(t,\kappa).$ (58)
Using Eq. (6) and (58), the expression (57) turns into
$\dfrac{1}{J(\kappa)}\dfrac{dJ(\kappa)}{d\kappa}$ $=$ $-i\int dt\Big[B(t,\kappa)\big(\dot{\zeta}(t,\kappa)-z(t,\kappa)\big)+\bar{\cal C}(t,\kappa)\,{\cal C}(t,\kappa)$ (59)
$+$ ${\dot{\bar{\cal C}}}(t,\kappa)\,\dot{\cal C}(t,\kappa)\Big].$
Further, we make an ansatz for $S_{1}[\phi(t,\kappa),\kappa]$ as follows:
$S_{1}[\phi(t,\kappa),\kappa]$ $=$ $i\int dt\Big[\xi_{1}(\kappa)\,B(t,\kappa)\,\dot{\zeta}(t,\kappa)+\xi_{2}(\kappa)\,B(t,\kappa)\,z(t,\kappa)+\xi_{3}\,(\kappa)\bar{\cal C}(t,\kappa)\,{\cal C}(t,\kappa)$ (60)
$+$ $\xi_{4}(\kappa)\,{\dot{\bar{\cal C}}}(t,\kappa)\,\dot{\cal C}(t,\kappa)\Big],$
where the coefficient $\xi_{i}\,(i=1,2,3,4)$ are arbitrary $\kappa$-dependent constant parameters. Now, differentiating $S_{1}[\phi(t,\kappa),\kappa]$ with respect to the parameter $\kappa$ by employing Eq. (6), we obtain
$\frac{dS_{1}}{d\kappa}$ $=$ $i\int dt\Big[\frac{d\xi_{1}(\kappa)}{d\kappa}\,B(t,\kappa)\,\dot{\zeta}(t,\kappa)+\frac{d\xi_{2}(\kappa)}{d\kappa}\,B(t,\kappa)\,z(t,\kappa)+\frac{d\xi_{3}(\kappa)}{d\kappa}\,{\cal\bar{C}}(t,\kappa)\,{\cal{C}}(t,\kappa)$ (61)
$+$ $\frac{d\xi_{4}(\kappa)}{d\kappa}\,{\dot{\bar{\cal{C}}}}(t,\kappa)\,\dot{\cal{C}}(t,\kappa)+\xi_{1}(\kappa)\,B(t,\kappa)\,\ddot{{\cal{C}}}(t,\kappa)\,\Theta^{\prime}+\xi_{2}(\kappa)\,B(t,\kappa)\,{\cal{C}}(t,\kappa)\,\Theta^{\prime}$
$+$ $\xi_{3}(\kappa)\,B(t,\kappa)\,{\cal{C}}(t,\kappa)\,\Theta^{\prime}+\xi_{4}(\kappa)\,\dot{B}(t,\kappa)\,\dot{\cal{C}}(t,\kappa)\,\Theta^{\prime}]\Big].$
Substituting Eqs. (59) and (61) into the necessary condition Eq. (56) leads to
$\int dt\Big[\Big(\frac{d\xi_{1}}{d\kappa}-i\Big)B\,\dot{\zeta}+\Big(\frac{d\xi_{2}}{d\kappa}+i\Big)B\,z+\Big(\frac{d\xi_{3}}{d\kappa}-i\Big){\cal\bar{C}}\,{\cal{C}}+\Big(\frac{d\xi_{4}}{d\kappa}-i\Big){\dot{\bar{\cal{C}}}}\,\dot{\cal{C}}$
$+\,\big(\xi_{2}+\xi_{3}\big)B\,{\cal{C}}\,\Theta^{\prime}+\big(\xi_{4}-\xi_{1}\big)\dot{B}\,\dot{\cal{C}}\,\Theta^{\prime}\Big]=0.$ (62)
Solving the above expression leads to a set of first-order linear differential equations as
$\frac{d\xi_{1}}{d\kappa}-i=0,\qquad\frac{d\xi_{2}}{d\kappa}+i=0,$
$\frac{d\xi_{3}}{d\kappa}-i=0,\qquad\frac{d\xi_{4}}{d\kappa}-i=0,$ (63)
and $\Theta^{\prime}$-dependent non-local terms vanish for
$\big(\xi_{2}+\xi_{3}\big)=\big(\xi_{4}-\xi_{1}\big)=0.$ (64)
The solutions of the differential equations under the initial condition $\xi_{i}(\kappa=0)=0$ is obtained as
$\xi_{1}=i\,\kappa,\qquad\xi_{2}=-i\,\kappa,\qquad\xi_{3}=i\,\kappa,\qquad\xi_{4}=i\,\kappa.$ (65)
Now, with these coefficients $\xi_{i}$, we can identify the functional $S_{1}[\phi(t,\kappa),\kappa]$ as follows:
$S_{1}[\phi(t,\kappa),\kappa]=\int dt\,\kappa\,\big(-B\,\dot{\zeta}+B\,z-\,\bar{\cal{C}}\,{\cal{C}}-\,{\dot{\bar{\cal{C}}}}\,\dot{\cal{C}}\big).$ (66)
Therefore, the total effective action, by adding this $S_{1}[\phi(t,\kappa),\kappa]$ to $S_{eff}$ in Eq. (49), can be written as follows
$S_{eff}+S_{1}[\phi(t,\kappa),\kappa]$ $=$ $\int dt\,\Big[\,\frac{1}{2}\,\dot{r}^{2}+\frac{1}{2}\,r^{2}\big(\dot{\theta}-g\zeta\big)^{2}+\frac{1}{2}\big(\dot{z}-\zeta\big)^{2}-V(r)$ (67)
$+$ $B\big(\dot{\zeta}-z+\frac{1}{2}B\big)-{\dot{\bar{\cal{C}}}}{\dot{\cal{C}}}-{\cal{\bar{C}}}{\cal{C}}+\kappa\big(-B\dot{\zeta}+Bz-\bar{\cal{C}}{\cal{C}}-{\dot{\bar{\cal{C}}}}\dot{\cal{C}}\Big].$
Interestingly, at $\kappa=0$, the above action simplified to
$S_{eff}+S_{1}[\phi(t,0),0]$ $=$ $\int dt\,\Big[\frac{1}{2}\,\dot{r}^{2}+\frac{1}{2}\,r^{2}\big(\dot{\theta}-g\zeta\big)^{2}+\frac{1}{2}\big(\dot{z}-\zeta\big)^{2}-V(r)$ (68)
$+$ $B\Big(\dot{\zeta}-z+\frac{1}{2}B\Big)-{\dot{\bar{\cal{C}}}}{\dot{\cal{C}}}-{\cal{\bar{C}}}{\cal{C}}\Big],$
which is same as the BFV-BRST gauge-fixed effective action in Eq. (49). However, at $\kappa=1$ the effective action in Eq. (67) reduces to the gauge-invariant classical action of the model as given below
$S_{eff}+S_{1}[\phi(t,1),1]$ $=$ $\int dt\,\Big[\frac{1}{2}\,\dot{r}^{2}+\frac{1}{2}\,r^{2}\big(\dot{\theta}-g\zeta\big)^{2}+\frac{1}{2}\big(\dot{z}-\zeta\big)^{2}-V(r)\Big].$ (69)
In the above, we eliminate the Nakanishi-Lautrup auxiliary variable $B$ and (anti-)ghost variables $(\bar{\cal{C}}){\cal{C}}$ using Gaussian integration. Thus, we demonstrate that, with a judicious choice of the parameter, FFBRST transformations connect the BFV-BRST gauge fixed action and the classical gauge-invariant action of the FLPR model.
At this juncture, it is worth mentioning that the finite field-dependent parameter in Cartesian coordinates is the same as that for polar coordinates. Therefore, the computation in the Cartesian coordinates can be performed in a similar fashion.